# Observing the exponential growth of the eigenmodes in the absence of coalescence for a non-Hermitian circuit with an unavoidable inductor dissipation

### v1

@@ Abstract
We investigate, both experimentally and theoretically, the eigenmodes of an electronic circuit in which gain and loss $RLC$ resonators are coupled through a capacitor. Due to the unavoidable magnetic loss in the inductors, we find that the eigenmode coalescence no longer emerges in contrast to the conventional non-Hermitian systems with the spontaneous $\cal{PT}$-symmetry breaking. In particular, we find a transition from the exponential decay to exponential growth in the amplitude of the periodic voltage oscillations of the resonators. The transition occurs near the exceptional points of the non-Hermitian circuit without considering the dissipations in inductors. We introduce a small resistor of three orders of magnitude smaller than that of the $RLC$ resonators to mimic the energy dissipation in inductors and numerically solve the equivalent non-Hermitian Schrödinger equation. The numerical results can well reproduce experimental observations. Our above findings unambiguously indicate that the exponential growth behavior beyond the exceptional points is robust against some unavoidable dissipative perturbations.
@@ II Model and main results
Figure 1: (a) Schematic of $\cal{PT}$-symmetric LC resonator. The negative resistance $-R$ in the gain resonator is provided by feedback from a voltage-doubling buffer. The two resonators are coupled by the capacitor $C_{0}$. (b) Double amplifier with $R_{2}=R_{3}$. (c) PCB board of the $\cal{PT}$-symmetric LC resonator. The buffer consists of high-impedance LF356 voltage followers, designed to isolate the influence of measurement instruments on the LC resonator.
@@
$V_{n}=R_{L}I_{n}^{L}+\frac{1}{L}\frac{dI_{n}^{L}}{dt}\;,$
where the superindices $R$, $L$, $C$, and $C_{0}$ indicate quantities associated with the resistor, inductor, capacitor and coupled capacitor, respectively. The subscripts denote the gain ($n=1$) and loss ($n=2$) resonators. Here, the current of the resistors and capacitors take the form $I_{n}^{R}=(-1)^{n}V_{n}/{R}$, $I_{1}^{C_{0}}=C_{0}({dV_{1}}/{dt}-{dV_{2}}/{dt})$, $I_{2}^{C_{0}}=-I_{1}^{C_{0}}$, and $I_{n}^{C}=C{dV_{n}}/{dt}$. For brevity, we introduce a set of dimensionless quantities $c={C}/{C_{0}}$, $k={R_{L}}/{R}$, $\gamma={R}^{-1}\sqrt{{L}/{C}}$, and $\gamma_{l}=R_{L}\sqrt{{C}/{L}}$. Dimensionless time is naturally defined as $\tau=\omega_{0}t$, where $\omega_{0}=1/\sqrt{LC}$ represents the resonant frequency. Straightforward calculations yield the relations
$\ddot{V}_{1}=$ $\dot{V}_{1}\frac{\gamma-(1+2c)\gamma_{l}}{c+1}+V_{1}\frac{k-1}{c+1}+\ddot{V}_{2}\frac{c}{c+1}+\dot{V}_{2}\frac{c\gamma_{l}}{c+1}\;,$ (2)
$\ddot{V}_{2}=$ $-\dot{V}_{2}\frac{\gamma+(1+2c)\gamma_{l}}{c+1}+V_{2}\frac{k+1}{c+1}+\ddot{V}_{1}\frac{c}{c+1}+\dot{V}_{1}\frac{c\gamma_{l}}{c+1}\;.$
@@
$M_{a}=\frac{1}{1+2c}\begin{bmatrix}{(1+c)(k-1)}&c(k+1)\\ {c(k-1)}&(1+c)(k+1)\end{bmatrix}\;,$ (5)
@@
We further perform experimental measurements of the eigenfrequencies of the coupled LC resonators. A comparison between $\omega$ and the numerical results is shown in Fig. 2. One can see that for $\gamma<\gamma_{c}$, the experimentally measured values of both $\omega_{r}$ and $\omega_{i}$ agree well with numerical simulations based on the model in Eq. (4). Note that we can extract only one imaginary component of the eigenfrequency spectrum because the difference between the two $\omega_{i}$ values is too small. For $\gamma>\gamma_{c}$, the eigenmode of the electronic circuit associated with a negative imaginary part $\omega_{i}$ decays exponentially. Therefore, only the eigenfrequency $\omega=\omega_{r}+i\omega_{i}$ with $\omega_{i}>0$ is experimentally accessible, and it is in perfect agreement with numerical results.
@@
We determine the eigenfrequency by applying a Fourier transform to the time evolution of the voltage $V_{j}$ at each node. Figure 3 shows the experimentally measured values of $V_{j}$ for different $\gamma$. One can see that for $\gamma<\gamma_{c}$, the $V_{j}$ periodically oscillates with time. The envelope of this oscillation is also periodic, while exhibiting an exponentially decaying amplitude [see Figs. 3(a) and (b)]. This reveals the existence of two real parts in the eigen-spectrum. The rate of the exponential decay of the envelop’s amplitude corresponds to the imaginary parts of the eigenvalues. For $\gamma\approx\gamma_{c}$, the amplitude of the envelope of $V_{j}$ is significantly larger compared to smaller values of $\gamma$. Interestingly, for $\gamma>\gamma_{c}$, the time evolution of $V_{j}$ exhibits periodic oscillations with a single frequency, and the amplitude increases exponentially with time. In this case, the oscillation frequency and the growth rate of the amplitude are $\omega_{r}$ and $\omega_{i}$, respectively. For comparison, we also employ the fourth-order Runge-Kutta algorithm to numerically solve Eq. (3) and obtain the time evolution of $V_{j}$. Our numerical results show perfect agreement with the experimentally measured values of $V_{j}$, confirming the validity of our model. This consistency is clear evidence that our introduction of $R_{L}$ to quantify energy dissipation in the coil $L$ is physically reasonable.
To implement $\mathcal{PT}$-symmetric circuits, gain components should be utilized to compensate for energy loss in inductors Schindler24APL ; Schindler12JPA . In this situation, for $\gamma<\gamma_{c}$, the voltage of the active $RLC$ resonators exhibits periodic oscillations due to energy gain. However, the unavoidable magnetic loss in inductors, which introduces small energy dissipation, leads to the exponential decay of voltage oscillations, corresponding to a negative value of $\omega_{i}$. Intrinsically, the eigenmodes of the voltage signal in our system exhibit a transition to exponential growth with time as the non-Hermitian parameter $\gamma$ crosses the EPs, indicating the robustness of exponential energy gain against perturbative dissipation.
@@ III Conclusion and discussions
The non-Hermiticity of electronic circuits generally arises from dissipative elements that introduce energy loss or gain, breaking energy conservation. In this work, we experimentally realize an electronic circuit where gain and loss $RLC$ resonators are coupled by a capacitor. The gain side is implemented using a voltage-doubling buffer, which provides feedback current with the relation $I_{out}=-V/R$, effectively achieving a negative resistance. Energy dissipation in the inductors is present in both the gain and loss $RLC$ resonators. Theoretically, we model this energy loss by introducing resistors $R_{L}$ in each inductor, thereby quantifying energy dissipation precisely. For $R_{L}=0$, the spontaneous $\cal{PT}$-symmetry breaking is controlled by a dimensionless parameter $\gamma$, with the EP occurring at $\gamma=\gamma_{c}$. We find that $V_{j}$ exhibits periodic oscillations over time, with its amplitude decaying exponentially for $\gamma<\gamma_{c}$ and increasing exponentially for $\gamma>\gamma_{c}$. Correspondingly, the imaginary part $\omega_{i}$ of the eigenfrequencies is negative for $\gamma<\gamma_{c}$ and positive for $\gamma>\gamma_{c}$. In addition, the two branches of $\omega_{i}$ diverge as $\gamma$ increases. The two real parts of the eigenfrequency converge as $\gamma$ increases but retain a gap for $\gamma>\gamma_{c}$. Therefore, the coalescence of eigenvalues is broken by the energy dissipation of inductors.

### v2

@@ Abstract
We investigate, both experimentally and theoretically, the eigenmodes of an electronic circuit in which gain and loss $RLC$ resonators are coupled through a capacitor. Due to the unavoidable magnetic loss in the inductors, we find that the eigenmode coalescence no longer emerges in contrast to the conventional non-Hermitian systems with the spontaneous $\cal{PT}$-symmetry breaking. In particular, we find a transition from the exponential decay to exponential growth in the amplitude of the periodic voltage oscillations of the resonators. The transition occurs near the exceptional points of the non-Hermitian circuit without considering the dissipations in inductors. We introduce a small resistor of two orders of magnitude smaller than that of the $RLC$ resonators to mimic the energy dissipation in inductors and numerically solve the equivalent non-Hermitian Schrödinger equation. The numerical results can well reproduce experimental observations. Our above findings unambiguously indicate that the exponential growth behavior beyond the exceptional points is robust against some unavoidable dissipative perturbations.
@@ II Model and main results
Figure 1: (a) Schematic of $\cal{PT}$-symmetric $RLC$ resonator. The negative resistance $-R$ in the gain resonator is provided by feedback from a voltage-doubling buffer. The two resonators are coupled by the capacitor $C_{0}$. (b) Double amplifier with $R_{2}=R_{3}$. (c) Printed circuit board of the $\cal{PT}$-symmetric $RLC$ resonator. The buffer consists of high-impedance LF356 voltage followers, designed to isolate the influence of measurement instruments on the $RLC$ resonator.
@@
$V_{n}=R_{L}I_{n}^{L}+L\frac{dI_{n}^{L}}{dt}\;,$
where the superindices $R$, $L$, $C$, and $C_{0}$ indicate quantities associated with the resistor, inductor, capacitor and coupled capacitor, respectively. The subscripts denote the gain ($n=1$) and loss ($n=2$) resonators. Here, the current of the resistors and capacitors take the form $I_{n}^{R}=(-1)^{n}V_{n}/{R}$, $I_{1}^{C_{0}}=C_{0}({dV_{1}}/{dt}-{dV_{2}}/{dt})$, $I_{2}^{C_{0}}=-I_{1}^{C_{0}}$, and $I_{n}^{C}=C{dV_{n}}/{dt}$. For brevity, we introduce a set of dimensionless quantities $c={C_{0}}/{C}$, $k={R_{L}}/{R}$, $\gamma={R}^{-1}\sqrt{{L}/{C}}$, and $\gamma_{l}=R_{L}\sqrt{{C}/{L}}$. Dimensionless time is naturally defined as $\tau=\omega_{0}t$, where $\omega_{0}=1/\sqrt{LC}$ represents the resonant frequency. Straightforward calculations yield the relations
$\ddot{V}_{1}=$ $\dot{V}_{1}\frac{\gamma-(1+c)\gamma_{l}}{c+1}+V_{1}\frac{k-1}{c+1}+\ddot{V}_{2}\frac{c}{c+1}+\dot{V}_{2}\frac{c\gamma_{l}}{c+1}\;,$ (2)
$\ddot{V}_{2}=$ $-\dot{V}_{2}\frac{\gamma+(1+c)\gamma_{l}}{c+1}-V_{2}\frac{k+1}{c+1}+\ddot{V}_{1}\frac{c}{c+1}+\dot{V}_{1}\frac{c\gamma_{l}}{c+1}\;.$
@@
$M_{a}=\frac{1}{1+2c}\begin{bmatrix}{(1+c)(k-1)}&-c(k+1)\\ {c(k-1)}&-(1+c)(k+1)\end{bmatrix}\;,$ (5)
@@
We further perform experimental measurements of the eigenfrequencies of the coupled $RLC$ resonators. A comparison between $\omega$ and the numerical results is shown in Fig. 2. One can see that for $\gamma<\gamma_{c}$, the experimentally measured values of both $\omega_{r}$ and $\omega_{i}$ agree well with numerical simulations based on the model in Eq. (4). Note that we can extract only one imaginary component of the eigenfrequency because the difference between the two $\omega_{i}$ values is too small. For $\gamma>\gamma_{c}$, the eigenmode of the electronic circuit associated with a negative imaginary part $\omega_{i}$ decays exponentially. Therefore, only the eigenfrequency with $\omega_{i}>0$ is experimentally accessible, and it is in perfect agreement with numerical results.
@@
We determine the eigenfrequency by applying a Fourier transform to the time evolution of the voltage $V_{j}$ at each node. Figure 3 shows the experimentally measured values of $V_{j}$ for different $\gamma$. One can see that for $\gamma<\gamma_{c}$, the $V_{j}$ periodically oscillates with time. The envelope of this oscillation is also periodic, while exhibiting an exponentially decaying amplitude [see Figs. 3(a) and (b)]. This reveals the existence of two real parts in the eigenfrequency. The rate of the exponential decay of the envelop’s amplitude corresponds to the imaginary parts of the eigenfrequency. For $\gamma\approx\gamma_{c}$, the amplitude of the envelope of $V_{j}$ is significantly larger compared to smaller values of $\gamma$. Interestingly, for $\gamma>\gamma_{c}$, the time evolution of $V_{j}$ exhibits periodic oscillations with a single frequency, and the amplitude increases exponentially with time. In this case, the oscillation frequency and the growth rate of the amplitude are $\omega_{r}$ and $\omega_{i}$, respectively. For comparison, we also employ the fourth-order Runge-Kutta algorithm to numerically solve Eq. (3) and obtain the time evolution of $V_{j}$. Our numerical results can well reproduce the experimentally measured values of $V_{j}$, confirming the validity of our model. This consistency is clear evidence that our introduction of $R_{L}$ to quantify energy dissipation in the coil $L$ is physically reasonable.
Non-Hermiticity is now widely recognized as an important extension of conventional quantum mechanics, providing a framework for describing open quantum systems. Notable phenomena, including the non-Hermitian many-body problem Xie19foP , biorthogonal dynamical phase transitions Jing24prl , and non-reciprocal adiabatic tunneling Chang25cpb , have demonstrated the fundamental significance of non-Hermitian physics. To implement $\mathcal{PT}$-symmetric circuits, gain components should be utilized to compensate for energy loss in inductors Schindler24APL ; Schindler12JPA . In this situation, for $\gamma<\gamma_{c}$, the voltage of the active $RLC$ resonators exhibits periodic oscillations due to energy gain. However, the unavoidable magnetic loss in inductors, which introduces small energy dissipation, leads to the exponential decay of voltage oscillations, corresponding to a negative value of $\omega_{i}$. Intrinsically, the eigenmodes of the voltage signal in our system exhibit a transition to exponential growth with time as the non-Hermitian parameter $\gamma$ crosses the EPs, indicating the robustness of exponential energy gain against perturbative dissipation.
@@ III Conclusion and discussions
The non-Hermiticity of electronic circuits generally arises from dissipative elements that introduce energy loss or gain, breaking energy conservation. In this work, we experimentally realize an electronic circuit where gain and loss $RLC$ resonators are coupled by a capacitor. The gain side is implemented using a voltage-doubling buffer, which provides feedback current with the relation $I_{out}=-V/R$, effectively achieving a negative resistance. Energy dissipation in the inductors is present in both the gain and loss $RLC$ resonators. Theoretically, we model this energy loss by introducing resistors $R_{L}$ in each inductor, thereby quantifying energy dissipation precisely. For $R_{L}=0$, the spontaneous $\cal{PT}$-symmetry breaking is controlled by a dimensionless parameter $\gamma$, with the EP occurring at $\gamma=\gamma_{c}$. We find that $V_{j}$ exhibits periodic oscillations over time, with its amplitude decaying exponentially for $\gamma<\gamma_{c}$ and increasing exponentially for $\gamma>\gamma_{c}$. Correspondingly, the imaginary part $\omega_{i}$ of the eigenfrequencies is negative for $\gamma<\gamma_{c}$ and positive for $\gamma>\gamma_{c}$. In addition, the two branches of $\omega_{i}$ diverge as $\gamma$ increases. The two real parts of the eigenfrequency converge as $\gamma$ increases but retain a gap for $\gamma>\gamma_{c}$. Therefore, the coalescence of eigenmodes is broken by the energy dissipation of inductors. Our finding that $\omega_{r}$ exhibits a gap for $\gamma>\gamma_{c}$ may represent a unique feature of non-Hermitian circuits, whose implications on sensing and wave manipulation warrant further investigation.